# SPARSE: Scattering Poles and Amplitudes from Radial Schrödinger Equations

### v2

@@ 3 Numerical solution
$r_{n}=nd$ (21)
with $d$ a small discretization step. Our discretized coordinate space begins at the origin $r_{0}=0$, has $M$ points in its interior that we refer to as nodes, and ends at a finite but otherwise arbitrarily large radius $r_{M+1}=R$. The numerical reduced radial wavefunction in channel $i$ is $u_{i}(r_{n})$. The numerical potential matrix element between the channels $i$ and $j$ is $V_{ij}(r_{n})$. The numerical second derivative of the reduced radial wavefunction is evaluated using the leading-order finite-difference method,
@@
where each $V_{i}^{\infty}$ is either a definite number or positive infinity. We identify $V_{i}^{\infty}$ with the energy threshold where the scattering channel $i$ opens up. The boundary condition for $u_{i}(R)$ depends on whether the energy $E$ in the Schrödinger equation is larger or smaller than the threshold $V_{i}^{\infty}$. If $E<V_{i}^{\infty}$, then a physical solution goes to zero for $r\to\infty$ and we impose $u_{i}(R)=0$. If otherwise $E>V_{i}^{\infty}$, a physical solution keeps oscillating indefinitely for $r\to\infty$. It is however a specific, finite number at the maximum distance $R$ in our truncated coordinate space. So, we impose $u_{i}(R)=b_{i}$, where $b_{i}$ is a real constant. We can conveniently impose the boundary condition for both cases $E<V_{i}^{\infty}$ and $E\geq V_{i}^{\infty}$ at once in the form
$u_{i}(R)=\theta(E-V_{i}^{\infty})b_{i},$ (25)
@@
$\sum_{j,m}\bigl[H_{ij,nm}-\delta_{ij}\delta_{nm}E\bigr]u_{j}(r_{m})\\ =\delta_{nM}\theta(E-V_{i}^{\infty})\frac{b_{i}}{2\mu_{i}d^{2}},$ (27)
@@
SPARSE reads this information from two comma-separated-values (CSV) files named channels.csv and potential.csv, which must be placed within the working directory.
@@
SPARSE reads the values from the two CSV files and stores them into two pandas DataFrame objects using the function read_csv from the pandas library [7]. From them, SPARSE calculates the Hamiltonian matrix elements $H_{ij,nm}$ from Equation (28) and writes them into a 2-dimensional NumPy array object [8] with elements $H_{\mu\nu}$. This reshaping amounts to applying an invertible map $(i,n)\mapsto\mu$ from pairs of indices $i=1,\dots,N$ and $n=1,\dots,M$ to a single index $\mu=1,\dots,NM$. SPARSE reads the matrix elements from Equation (28) and writes them in the NumPy array following a C-style index order, which corresponds to setting
@@
Therefore the $H_{\mu\nu}$ form the elements of a banded matrix where all the nonzero elements are contained within the main diagonal, the first $N$ upper diagonals, and the first $N$ lower diagonals.
@@
$B_{\mu}(E)=\delta_{nM}\theta(E-V_{i}^{\infty})\frac{b_{i}}{2\mu_{i}d^{2}}.$ (32)
Note that if $E<V_{i}^{\infty}$ for all $i$, which corresponds to a bound state, then $B_{\mu}(E)=0$ for all $\mu$ and Equation (31) becomes an eigenvalue problem for the Hamiltonian matrix. Even though SPARSE is primarily designed for calculating scattering states, it also contains a handy bound-state calculator which solves Equation (31) with $B_{\mu}(E)=0$ using the function eigsh from the sparse linear algebra module of SciPy [9].
Finally, a note about measuring units. SPARSE uses natural units, in which $\hbar=c=1$ and therefore $\text{Time}]=[\text{Length}]$, $\text{Mass}]=[\text{Energy}]$, and $\text{Length}]=[\text{Energy}]^{-1}$. There is only one independent dimension, typically chosen between $\text{Length}]$ in units of fm (femtometers) and $\text{Energy}]$ in units of eV (electron Volts). The conversion between length-based and energy-based units is easily achieved by means of the formula $\hbar c=1=197.327$ MeV fm. The SPARSE algorithm is agnostic to the user’s choice of units, but it assumes that the dimension of the inputs are $r_{n}]=[\text{Length}]$ and $V_{ij}(r_{n})]=[\mu_{i}]=[V_{i}^{\infty}]=[\text{Energy}]$.
@@ 4 Calculating the $K$-matrix
For a given energy $E$, the possible values of the constant in Equation (32) constitute a vector space of dimension $O$, where $O$ is the number of open channels $i$ whose threshold is less than $E$, $V_{i}^{\infty}\leq E$ for $i=1,\dots,O$. We can easily construct a basis for this vector space by substituting $b_{i}$ inside Equation (32) with
$b_{ij}=\delta_{ij},$ (33)
@@
with $I$ the identity matrix. $A$ is a banded matrix which SPARSE calculates by shifting the main diagonal of the numerical Hamiltonian by E-E in the matrix diagonal ordered form. With these definitions, we rewrite Equation (31) as a matrix equation
@@
where $u$ is a $(NM)\times O$ matrix whose columns are the linearly independent solutions to Equation (31) with the same energy $E$.
For any energy $E$, SPARSE constructs the matrices $A$ and $B$ using Equations (32)-(34) and solves Equation (35) for the matrix $u$ using the function solve_banded from the linear algebra module of SciPy [9]. Then, SPARSE reshapes the matrix $u$ into a set of numerical reduced wavefunctions using Equation (29): $\mu\to(i,n)$, $u_{\mu j}\to u_{ij}(r_{n})$. Next, we use these numerical wavefunctions to calculate the reactance matrix $K$ for the scattering process. We will disregard from now on wavefunction components for confining channels $i$ with $V_{i}^{\infty}=\infty$, which are irrelevant to the scattering states at large $r$.
@@
Next, we review the analytic wavefunctions of the scattering states at large $r$. For $r\gg r_{V}$, Equation (20) reduces to a set of decoupled Schrödinger equations for free particles,
$\biggl[-\frac{1}{2\mu_{i}}\frac{\mathrm{d}^{2}\hphantom{r}}{\mathrm{d}r^{2}}+\frac{l_{i}(l_{i}+1)}{2\mu r^{2}}+V_{i}^{\infty}-E\biggr]u_{i}(r)=0,$ (37)
for every channel $i$ with a finite threshold $V_{i}^{\infty}$. For a closed channel $i$ with $V_{i}^{\infty}>E$ ($i=O+1,\dots,N$), a physical solution to Equation (37) must approach zero at large distances. For an open channel $i$ with $V_{i}^{\infty}\leq E$ ($i=1,\dots,O$), Equation (37) has two linearly independent physical solutions. Using real boundary conditions and wavefunction normalization by energy (see Reference [10] for alternative conventions), the two solutions are
@@
where the relation $\simeq$” stands for equality at large $r$, the first index $i=1,\dots,O$ labels open channels, the second index $j=1,\dots,O$ labels linearly independent solutions with the same energy, and $K_{ij}$ are $K$-matrix elements (see Section 11.1.2 of Reference [11]).
@@
Therefore, we can obtain the numerical values of $X_{ij}$ and $Y_{ij}$ by evaluating the integrals $\mathcal{S}_{ij}$ and $\mathcal{C}_{ij}$ in Equation (46) using the numerical wavefunctions $u_{ij}(r_{n})$ and Simpson’s rule. SPARSE uses the simpson function from the integration module of SciPy [9]. As long as the momenta $p_{i}$ satisfy $p_{i}\ll d^{-1}$, the integrands of $\mathcal{S}_{ij}$ and $\mathcal{C}_{ij}$ are slowly oscillating functions with respect to the discretization step $d$. Then Simpson’s rule provides an accurate numerical estimate of the integrals $\mathcal{S}_{ij}$ and $\mathcal{C}_{ij}$ in Equation (46).
@@ 5 Scattering poles and amplitudes
(see Section 14-e of Reference [12]). We see from Equations (54) and (55) that unitarity ($S^{\dagger}=S^{-1}$) and time-reversal symmetry ($S^{\mathsf{T}}=S$) require that the $K$-matrix is real and symmetric. The $K$-matrix calculated using Equations (44), (46), and (51) is real by definition. It is only approximately symmetric within numerical errors. Let us call this approximately symmetric matrix $\tilde{K}$. SPARSE calculates the degree of asymmetry as
@@
Resonances may appear as poles of the $K$-matrix on the real axis of the physical Riemann sheet. For energies near the pole, the $K$-matrix elements can be expressed as
@@
SPARSE calculates scattering poles by interpolating between numerical $K$-matrix elements obtained at a set of discrete energies. The interpolation is performed using the AAA algorithm for rational polynomial approximation [13], which yields the pole energies $M$ and the residues of the $K$-matrix elements at the poles,
$R_{ij}=\operatorname{Res}_{M}(K_{ij}).$ (60)
Not all poles of the $K$-matrix correspond to resonances. SPARSE identifies $K$-matrix poles associated to resonances by comparing the residues calculated from the AAA algorithm with the analytic pole term in Equation (58). If the $K$-matrix pole indeed corresponds to a resonance, then Equations (58) and (60) require
@@
where the sign function on the right side is necessary to fix the relative signs between the couplings so that they agree with Equation (61). Note that the couplings are defined up to an overall sign, that is, the $K$-matrix in Equation (58) is invariant if we substitute $g_{i}\to-g_{i}$ for every $i$. In writing Equation (64), we have chosen the convention $g_{i}\geq 0$ for $i=1$.
@@
and using Equation (53). The corresponding partial-wave cross-section, $\sigma_{i\leftarrow j}$, is proportional to the square modulus of the scattering amplitudes,
@@
where $J$ is the total angular momentum and $J_{1,j}$ and $J_{2,j}$ are the spins of the particles 1 and 2 in the initial channel $j$. The total unpolarized cross section can be obtained by summing over the spins $s_{i}$, $s_{j}$, the orbital angular momenta $l_{i}$, $l_{j}$, and the total angular momentum $J$.
@@
If the resonance is not isolated, then the parameters $M$, $\Gamma$, and $g_{i}$ obtained from the $K$-matrix pole can differ significantly from those of the $T$-matrix pole. Nonetheless, the $K$-matrix pole parameters are physically relevant as they can be related to $K$-matrix fits to experimental data. In those cases where a distinction is necessary, $M$, $\Gamma$, and $g_{i}$ may be referred to as the nominal mass, width, and couplings of the resonance.
@@ 6 Showcase
Table 1: Integer label $i$, orbital angular momentum $l$, reduced mass $\mu$, and threshold $V^{\infty}$ for the channels in the SPARSE showcase example. Units are arbitrary.
$i$ $l$ $\mu$ $V^{\infty}$
@@
Next, we examine an application of the SPARSE algorithm to a relatively simple system with two coupled channels. Units are arbitrary. The two channels are described in Table 1. The potential matrix is
@@
$V_{11}(r)=\biggl(\frac{0.05}{r}-100\biggr)\exp\bigl(-(100r)^{2}\bigr).$ (69)
The second diagonal element $V_{22}(r)$ is the same potential well but shifted by +100+100,
$V_{22}(r)=V_{11}(r)+100.$ (70)
@@
$V_{12}(r)=50(100r)^{2}\exp\bigl(-(100r)^{2}\bigr).$ (71)
The coordinate space of $r$ is a grid of points in the interval $[0,1]$ with a discretization step $d=10^{-6}$. The potential $V_{11}(r)$ has been engineered so that it has a $P$-wave (that is, $l=1$) bound state with energy somewhat below the first threshold, $V^{\infty}_{1}=0$. The potential $V_{22}(r)$ has been engineered so that in isolation it would have an $S$-wave (that is, $l=0$) bound state with energy somewhat below the second threshold, $V^{\infty}_{2}=100$. The coupling potential $V_{12}(r)$ has been engineered so that the would-be bound state in the second channel appears instead as a narrow scattering resonance in the first channel. These potential matrix elements are plotted in Figure 1.
Figure 1: Potential matrix elements for the SPARSE showcase example. Axes have arbitrary units.
We use SPARSE to calculate the $K$-matrix for a grid of energies in the interval $[0,80]$ with a constant spacing of 0.05. The $K$-matrix has only one element, since there is only one open channel in this energy region. We plot the $K$-matrix as a function of the energy in the left panel of Figure 2. We also calculate the square modulus of the $T$-matrix element and plot it in the right panel of Figure 2.
Figure 2: $K$-matrix (left) and square modulus of the $T$-matrix (right) in the SPARSE showcase example. The horizontal axes have arbitrary units.
The scattering amplitude looks complicated because it results from the overlap of two different line shapes. The smooth enhancement is due to the bound state in channel 1, which has an energy of -12.5. The sharp oscillation of the amplitude in between 60 and 80 is due to a resonance. That oscillation in $\lvert T\rvert^{2}$ matches a pole of the $K$-matrix with mass $M=67.09$ and width $\Gamma=0.66$.
The SPARSE repository contains a Python script that prepares the CSV files channels.csv and potential.csv used for this example. It also contains a handy IPython notebook which reproduces the results of this section using the SPARSE Application Programming Interface. The reader is invited to go through the IPython notebook as a tutorial for using SPARSE.

### v3

@@ 3 Numerical solution
$r_{n}=nd,$ (26)
with $d$ a small discretization step. Our discretized coordinate space begins at the origin $r_{0}=0$, has $M$ points in its interior that we refer to as nodes, and ends at a finite but otherwise arbitrarily large radius $r_{M+1}=R$. The numerical reduced radial wavefunction in the channel $i$ is $u_{i}(r_{n})$. The numerical potential matrix element coupling the channel $j$ to the channel $i$ is $V_{ij}(r_{n})$. The numerical second derivative of the reduced radial wavefunction is evaluated using the leading-order finite-difference method,
@@
where each $V_{i}^{\infty}$ is either a definite number or positive infinity. We identify $V_{i}^{\infty}$ with the energy threshold where the scattering channel $i$ opens up. The boundary condition for $u_{i}(R)$ depends on whether the energy $E$ in the Schrödinger equation is larger or smaller than the threshold $V_{i}^{\infty}$. If $E<V_{i}^{\infty}$, then a physical solution goes to zero for $r\to\infty$ and we impose $u_{i}(R)=0$. If otherwise $E>V_{i}^{\infty}$, a physical solution keeps oscillating indefinitely for $r\to\infty$. It is however a specific, finite number at the maximum distance $R$ in our truncated coordinate space. So, we impose $u_{i}(R)=U_{i}$, where $U_{i}$ is a real constant. We can conveniently impose the boundary condition for both cases $E<V_{i}^{\infty}$ and $E\geq V_{i}^{\infty}$ at once in the form
$u_{i}(R)=\theta(E-V_{i}^{\infty})U_{i},$ (30)
@@
$\sum_{j,m}\bigl[H_{ij,nm}-\delta_{ij}\delta_{nm}E\bigr]u_{j}(r_{m})\\ =\delta_{nM}\theta(E-V_{i}^{\infty})\frac{U_{i}}{2\mu_{i}d^{2}},$ (32)
@@
SPARSE reads this information from two comma-separated-values (CSV) files named channels.csv and potential.csv, which must be placed within the working directory. Note that all the required spin information is encoded in the potential matrix elements through Equation (22). The final input, the total energy $E$, will be specified by the user as an argument of the functions defined in the SPARSE module.
@@
SPARSE reads the values from the two CSV files and stores them into two pandas DataFrame objects using the function read_csv from the pandas library [undeff]. From them, SPARSE calculates the Hamiltonian matrix elements $H_{ij,nm}$ from Equation (33) and writes them into a 2-dimensional NumPy array object [undefg] with elements $H_{\mu\nu}$. This reshaping amounts to applying an invertible map $(i,n)\mapsto\mu$ from pairs of indices $i=1,\dots,N$ and $n=1,\dots,M$ to a single index $\mu=1,\dots,NM$. SPARSE reads the matrix elements from Equation (33) and writes them in the NumPy array following a C-style index order, which corresponds to setting
@@
Therefore the $H_{\mu\nu}$ form a banded matrix where all the nonzero elements are contained within the main diagonal, the first $N$ upper diagonals, and the first $N$ lower diagonals.
@@
$B_{\mu}(E)=\delta_{nM}\theta(E-V_{i}^{\infty})\frac{U_{i}}{2\mu_{i}d^{2}}.$ (37)
Note that if $E<V_{i}^{\infty}$ for all $i$, which corresponds to a bound state, then $B_{\mu}(E)=0$ for all $\mu$ and Equation (36) becomes an eigenvalue problem for the Hamiltonian matrix. Even though SPARSE is primarily designed for calculating scattering states, it also contains a handy bound-state calculator which solves Equation (36) with $B_{\mu}(E)=0$ using the function eigsh from the sparse linear algebra module of SciPy [undefh].
Finally, a note about measuring units. SPARSE uses natural units, in which $\hbar=c=1$ and therefore $\text{Time}]=[\text{Length}]$, $\text{Mass}]=[\text{Energy}]$, and $\text{Length}]=[\text{Energy}]^{-1}$. There is only one independent dimension, typically chosen between $\text{Length}]$ in units of fm (femtometers) and $\text{Energy}]$ in units of eV (electron Volts). The conversion between length-based and energy-based units is easily achieved by means of the formula $\hbar c=1=197.327$ MeV fm. The SPARSE algorithm is agnostic to the user’s choice between length-based or energy-based units, but it assumes that the dimensions of the inputs are consistent with $r_{n}]=[\text{Length}]$ and $V_{ij}(r_{n})]=[\mu_{i}]=[V_{i}^{\infty}]=[\text{Energy}]$.
@@ 4 Calculating the $K$-matrix
For a given energy $E$, the possible values of the constant in Equation (37) constitute a vector space of dimension $O$, where $O$ is the number of open channels $i$ whose threshold is less than $E$, $V_{i}^{\infty}\leq E$ for $i=1,\dots,O$. We can easily construct a basis for this vector space by substituting $U_{i}$ inside Equation (37) with
$U_{ij}=\delta_{ij},$ (38)
@@
with $I$ the identity matrix. In practice, $A$ is a banded matrix which SPARSE calculates by applying an offset of E-E to the main diagonal of the numerical Hamiltonian in the matrix diagonal ordered form. With these definitions, we rewrite Equation (36) as a matrix equation
@@
where $u$ is a $(NM)\times O$ matrix whose columns form a complete set of linearly independent solutions to Equation (36) with the same energy $E$.
For any energy $E$, SPARSE constructs the matrices $A$ and $B$ using Equations (37)-(39) and solves Equation (40) for the matrix $u$ using the function solve_banded from the linear algebra module of SciPy [undefh]. Then, SPARSE reshapes the matrix $u$ into a set of numerical reduced wavefunctions using Equation (34): $\mu\to(i,n)$, $u_{\mu j}\to u_{ij}(r_{n})$. Next, we use these numerical wavefunctions to calculate the reactance matrix $K$ for the scattering process. We will disregard from now on wavefunction components for confining channels $i$ with $V_{i}^{\infty}=\infty$, which are irrelevant to the scattering states at large $r$.
@@
Next, we review the analytic wavefunctions of the scattering states at large $r$. For $r\gg r_{V}$, the system in Equation (25) reduces to a set of decoupled Schrödinger equations for free particles,
$\biggl[-\frac{1}{2\mu_{i}}\frac{\mathrm{d}^{2}\hphantom{r}}{\mathrm{d}r^{2}}+\frac{l_{i}(l_{i}+1)}{2\mu r^{2}}+V_{i}^{\infty}-E\biggr]u_{i}(r)=0.$ (42)
For a closed channel $i$ with $V_{i}^{\infty}>E$ ($i=O+1,\dots,N$), a physical solution to Equation (42) must approach zero at large distances. For an open channel $i$ with $V_{i}^{\infty}\leq E$ ($i=1,\dots,O$), Equation (42) has two linearly independent physical solutions. Using real boundary conditions and wavefunction normalization by energy (see Reference [undefi] for alternative conventions), the two solutions are
@@
where the relation $\simeq$” stands for equality at large $r$, the first index $i=1,\dots,O$ labels open channels, the second index $j=1,\dots,O$ labels linearly independent solutions with the same energy, and $K_{ij}$ are $K$-matrix elements (see Section 11.1.2 of Reference [undefj]).
@@
Therefore, we can obtain the numerical values of $X_{ij}$ and $Y_{ij}$ by evaluating the integrals $\mathcal{S}_{ij}$ and $\mathcal{C}_{ij}$ in Equation (51) using the numerical wavefunctions $u_{ij}(r_{n})$ and Simpson’s rule. SPARSE uses the simpson function from the integration module of SciPy [undefh]. As long as the momenta $p_{i}$ satisfy $p_{i}\ll d^{-1}$, the integrands of $\mathcal{S}_{ij}$ and $\mathcal{C}_{ij}$ are slowly oscillating functions with respect to the discretization step $d$. Then Simpson’s rule provides an accurate numerical estimate of the integrals $\mathcal{S}_{ij}$ and $\mathcal{C}_{ij}$ in Equation (51).
@@ 5 Scattering poles and amplitudes
(see Section 14-e of Reference [undefk]). The poles of the $K$-matrix are located on the real axis of the physical Riemann sheet. We see from Equations (59) and (60) that unitarity ($S^{\dagger}=S^{-1}$) and time-reversal symmetry ($S^{\mathsf{T}}=S$) require that the $K$-matrix is real and symmetric. The $K$-matrix calculated using Equations (49), (51), and (56) is real by definition. It is only approximately symmetric within numerical errors. Let us call this approximately symmetric matrix $\tilde{K}$. SPARSE calculates the degree of asymmetry as
@@
Resonances may appear as poles of the $K$-matrix on the real axis of the physical Riemann sheet. For energies near the pole, the $K$-matrix elements can be expressed as [undefl]
@@
SPARSE calculates scattering poles by interpolating between numerical $K$-matrix elements obtained at a set of discrete energies. The interpolation is performed using the AAA algorithm for rational polynomial approximation [undefm]. For a function $F$ evaluated at a discret set of energies, the interpolating rational function is represented in the barycentric form,
$\frac{N(E)}{D(E)}=\frac{\sum_{n=1}^{m}w_{n}F_{n}/(E-E_{n})}{\sum_{n=1}^{m}w_{n}/(E-E_{n})},$ (65)
where $m$ is the number of iterations of the AAA algorithm, $E_{n}$ are support points selected from the provided set of discrete energies, $F_{n}=F(E_{n})$ are the corresponding values of $F$, and $w_{n}$ are weights. The polynomials $N(E)$ and $D(E)$ can be obtained respectively by multiplying the numerator and the denominator of the right side of Equation (65) by $\prod_{n=1}^{m}(E-E_{n})$. We see then that $N(E)$ and $D(E)$ are polynomials of degree m1m-1. The interpolation in Equation (65) is exact only at the support points,
$F(E_{n})=\frac{N(E_{n})}{D(E_{n})},$ (66)
and it is approximate elsewhere. The number of iterations $m$ can be fixed by requiring the accuracy of the interpolation to be within a certain tolerance. Alternatively, one can fix $m$ to the number of poles plus one. From Equation (65), we see that the zeros, poles, and residues of the interpolating function can be computed in terms of the support points $E_{n}$, the corresponding function values $F_{n}$, and the weights $w_{n}$. A more detailed description of the AAA algorithm for rational polynomial approximation, which goes beyond the scope of this paper, can be found in Reference [undefm].
SPARSE estimates the number of poles of a $K$-matrix element $K_{ij}$ evaluated at a set of discrete energies by counting the number of instances in which $K_{ij}$ is large and changes sign between nearest-neighbor energies. Then, it applies the AAA function from the interpolation module of SciPy [undefh] with a number of iterations fixed to the number of poles plus one. Repeating this procedure for every element of the $K$-matrtix yields the energy $M$ and the residue matrix
$R_{ij}=\operatorname{Res}_{M}(K_{ij})$ (67)
of each pole. [^7] SPARSE checks that the pole energies are the same for all $K_{ij}$ up to a small tolerance. Not all poles of the $K$-matrix correspond to resonances. SPARSE identifies $K$-matrix poles associated to resonances by comparing the residues calculated from the AAA algorithm with the analytic pole term in Equation (63). If the $K$-matrix pole indeed corresponds to a resonance, then Equations (63) and (67) require
@@
where the sign function on the right side is necessary to fix the relative signs between the couplings so that they agree with Equation (68). Note that the couplings are defined up to an overall sign, that is, the $K$-matrix in Equation (63) stays the same if we substitute $g_{i}\to-g_{i}$ for every $i$. In writing Equation (71), we have chosen the convention $g_{i}\geq 0$ for $i=1$.
@@
and using Equation (58). The partial-wave cross-section, $\sigma_{i\leftarrow j}$, is proportional to the square modulus of the scattering amplitude,
@@
where $J$ is the total angular momentum and $J_{1,j}$ and $J_{2,j}$ are the spins of the interacting particles in the initial channel $j$. The total unpolarized cross section can be obtained by summing over the spins $s_{i}$, $s_{j}$, the orbital angular momenta $l_{i}$, $l_{j}$, and the total angular momentum $J$.
@@
If the resonance is not isolated, that is, if the “nonresonant” background under the pole in Eq. (63) is not small enough for energies $\lvert E-M\rvert\lesssim\Gamma$, then the parameters $M$, $\Gamma$, and $g_{i}$ obtained from the $K$-matrix pole can differ significantly from those of the $T$-matrix pole. Nonetheless, the $K$-matrix pole parameters are physically relevant as they can be extracted from $K$-matrix fits to experimental data. In those cases where a distinction is necessary, $M$, $\Gamma$, and $g_{i}$ are sometimes referred to as the nominal mass, width, and couplings of the resonance.
@@ 6 Showcase
Table 1: Integer label $i$, orbital angular momentum $l$, reduced mass $\mu$, and threshold $V^{\infty}$ for the channels in the SPARSE showcase example.
$i$ $l$ $\mu$ [MeV] $V^{\infty}$ [MeV]
@@
Next, we examine an application of the SPARSE algorithm to a relatively simple system with two coupled channels. We use energy-based natural units where all physical quantities are measured in MeV. The two channels are described in Table 1. The potential matrix is
@@
$V_{11}(r)=\biggl(\frac{\kappa}{r}-E_{0}\biggr)\exp\bigl(-r^{2}/r_{0}^{2}\bigr),$ (76)
where $\kappa=0.05$, $E_{0}=100$ MeV, and $r_{0}=10^{-2}$ MeV-1. The second diagonal element $V_{22}(r)$ is the same potential well but shifted upwards by $E_{0}$,
$V_{22}(r)=V_{11}(r)+E_{0}.$ (77)
@@
$V_{12}(r)=Gr^{2}/r_{0}^{2}\exp\bigl(-r^{2}/r_{0}^{2}\bigr),$ (78)
where $G=50$ MeV. The coordinate space of $r$ is a grid of points in the interval $[0,1]$ MeV-1 with a discretization step $d=10^{-6}$ MeV-1. The potential $V_{11}(r)$ has been engineered so that it has a $P$-wave (that is, $l=1$) bound state with energy somewhat below the first threshold, $V^{\infty}_{1}=0$. The potential $V_{22}(r)$ has been engineered so that in isolation it would have an $S$-wave (that is, $l=0$) bound state with energy somewhat below the second threshold, $V^{\infty}_{2}=E_{0}$. The coupling potential $V_{12}(r)$ has been engineered so that the would-be bound state in the second channel appears instead as a narrow scattering resonance in the first channel. These potential matrix elements are plotted in Figure 1.
Figure 1: Potential matrix elements for the SPARSE showcase example.
We use SPARSE to calculate the $K$-matrix for a grid of energies in the interval $[0,80]$ MeV with a constant spacing of 0.05 MeV. The $K$-matrix has only one element, since there is only one open channel in this energy region. We plot the $K$-matrix as a function of the energy in the left panel of Figure 2. We also calculate the square modulus of the $T$-matrix element and plot it in the right panel of Figure 2.
Figure 2: $K$-matrix (left) and square modulus of the $T$-matrix (right) in the SPARSE showcase example.
The scattering amplitude looks complicated because it results from the overlap of two different line shapes. The smooth enhancement is due to the bound state in channel 1, which has an energy of -12.5 MeV. The sharp oscillation of the amplitude in between 60 MeV and 80 MeV is due to a resonance. That oscillation in $\lvert T\rvert^{2}$ matches a pole of the $K$-matrix with mass $M=67.09$ MeV and width $\Gamma=0.66$ MeV.
The SPARSE repository contains a Python script that prepares the CSV files channels.csv and potential.csv used for this example. It also contains a handy IPython notebook which reproduces the results of this section using the SPARSE Application Programming Interface. The reader is invited to execute the IPython notebook as a tutorial for using SPARSE.